solve
y=2x-1
2x-y-1=0
---
So far i have done
2x-y-1=0
+y
2x-1=y
But it's the same equation as the top one. Could you please tell me how to solve this system?

Found 2 solutions by mananth, ewatrrr:
Answer by mananth(16946)

y=2x-1
2x-y-1=0

y=2x-1
y=2x-1
Both the equations represent the same line.
so the system has innumerable solutions.

Answer by ewatrrr(24785)

 
Hi,
solve
y=2x-1
2x-y-1=0
2x-y = 1
2x-y = 1 ||Yes! Same Equations. Dependent System. Infinite amount of solutions
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions. ***